OpenStudy (anonymous):

Find f^-1(x) f(x) 3 square-root of x-3

geerky42 (geerky42):

\(f(x) = 3\sqrt{x-3}\) ?

geerky42 (geerky42):

Or \(\sqrt[3]{x-3}\)?

OpenStudy (anonymous):

the second one

geerky42 (geerky42):

Switch x and y to each other then isolate y. Like this: \[x = \sqrt[3]{y-3}\]Now cube both sides.\[x^3 = y-3\]\[\Large y = f^{-1}(x) = \boxed{x^3 + 3}\]

OpenStudy (anonymous):

Thank you.

geerky42 (geerky42):

You're welcome.